the primeter of a triangle is 144. What are the consecutive integers for reach side
Answer by stanbon(75887)

the primeter of a triangle is 144. What are the consecutive integers for reach side
------------
x + (x+1) + (x+2) = 144
3x + 3 = 144
3x = 141
x = 47
x+1 = 48
x+2 = 49
========================
Cheers,
Stan H.